Find the surface area of a cylinder inscribed in a cube with the given sides.
8 cm

Answer by mananth(16949)

Let the cube side be a
radius of cylinder =a/2
height of cylinder = a
surface area of cylinder = 2pir(r+h)
=2 * pi*a/2*((a/2)+a)
=2*pi*(a/2)*(3a/2)
=(3/2)*pi*a^2